What is the [H3O+] in a 0.63 M acetic (HC2H3O2) solution that is 5.8% ionized?


1
Expert's answer
2021-05-27T07:35:06-0400

Strong electrolytes dissociate completely in the aqueous medium and have a high dissociation constant while the weak electrolytes dissolve partially in the aqueous medium and have a low ionization constant.

From the ionization constant and the degree of the ionization constant, we can calculate the value of the concentration of hydrogen ions and hydroxide ions in the solution.


Acetic acid is dissociated as

CH3COOH"\implies" CH3COO- + H+

"C(1-\alpha)" "C\alpha" "C\alpha"

Ka"=\frac{C(1-\alpha)}{C\alpha.C\alpha}"

Ka"=\frac{1-\alpha}{C\alpha^2}"


Given "\alpha=0.058"

C="0.63M"

Therefore,

[H+]=[H3O+]="C\alpha" ="0.63\u00d70.058"


="0.036M"